# Unveiling the Unexplored Decay Mode of a Light Charged Higgs Boson to an Off-Shell Top Quark and a Bottom Quark

### v1

@@ IV $H^{\pm}\rightarrow t^{*}b$ at a Multi-TeV Muon Collider
The Drell-Yan process exhibits nearly constant cross sections for different $M_{H^{\pm}}$ values at a given $\sqrt{s}$: approximately 3 fb at $\sqrt{s}=3{\;{\rm TeV}}$ and 0.3 fb at $\sqrt{s}=10{\;{\rm TeV}}$. This consistency is due to the relatively light $M_{H^{\pm}}$ compared to the multi-TeV collision energy. The marked decrease in cross sections from $\sqrt{s}=3{\;{\rm TeV}}$ to $\sqrt{s}=10{\;{\rm TeV}}$ illustrates the typical behavior of Drell-Yan cross sections, which are inversely proportional to the square of the beam energy. In contrast, the cross sections for the second and third processes moderately increase with increasing $\sqrt{s}$. This opposing trend is characteristic of VBS processes, whose cross sections increase according to $\log^{2}({s}/{m_{V}^{2}})$ Han:2021udl .
@@
Next, we present the signal-to-background analysis at the detector level. We conducted showering with Pythia version 8.307. A rapid detector simulation was performed using Delphes version 3.5.0, utilizing the delphes_card_MuonColliderDet.tcl card. [^4] It is accessible at https://github.com/delphes/delphes/blob/master/cards/delphes_card_MuonColliderDet.tcl the MuC Delphes card accommodates slightly different values for the $\tau$ tagging and mistagging rates compared to those for the HL-LHC: for $p_{T}\geq 10{\;{\rm GeV}}$, $P_{\tau\to\tau}\simeq 80\%$, $P_{e\rightarrow\tau}\simeq 0.1\%$, and $P_{j\to\tau}\simeq 2\%$.
For jet clustering, we employed the exclusive Valencia algorithm Boronat:2014hva ; Boronat:2016tgd , implemented in FastJet. This algorithm is particularly well-suited for high-energy lepton colliders due to its adept handling of initial state radiation and beam-induced backgrounds by incorporating beam jets. We chose the exclusive clustering setting for two primary reasons. First, it is more efficient for signals with a predefined number of jets, such as our signal $H^{+}H^{-}\rightarrow t^{*}b\tau\nu$, which requires five jets. The exclusive algorithm terminates jet clustering when the jet count matches the specified number of jets. Second, the inclusive algorithm is often ineffective for our signal jets since it can easily miss very close jets due to its requirement for a minimum distance between any two jets. Our signal jets, originating from the decay of the same parent particle $H^{\pm}$ produced at the multi-TeV MuC, tend to have a small jet radius and thus can fail to be selected by the inclusive algorithm. To overcome this, we implemented the exclusive Valencia Algorithm with a jet radius of $R=0.2$ and $\alpha=\beta=1$.
@@
To discuss the potential SM backgrounds for the final state $bbjj\tau\nu$ at the MuC, we first note that producing the $\tau\nu$ system requires the presence of a $W^{\pm}$ boson. Since the $W^{\pm}$ boson must be pair-produced at the MuC, the two light jets in the final state originate from the decay of the other $W^{\pm}$ boson. Considering the possibility of light quark jets being misidentified as $b$ jets, we identify two primary background processes: $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}b\bar{b}$ and $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}jj$. It is worth noting that the $W^{+}W^{-}b\bar{b}$ process encompasses top quark pair production, whereas $W^{+}W^{-}jj$ includes the contribution from $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}Z$. After applying $b$-jet mistagging rates, we found that the contribution from the $W^{+}W^{-}jj$ process is negligible. Therefore, we identify $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}b\bar{b}$ as the main background for the $bbjj\tau\nu$ final state.
@@
Brief comments on BIB are in order here. Beam-induced backgrounds arise from the decay of muons in the beam, which produces electrons and positrons. As these interact with machine components, various secondary particles such as photons, electron-positron pairs, hadrons, and neutrinos are generated. Many of these secondary particles result from forward scattering, causing most BIB particles to be directed along the beam and to have low energies, typically below 1 GeV Bartosik:2024ulr . Given that our final state requires five hard jets in the central region, BIB are of negligible concern. Thus, we do not consider BIB a significant background for our analysis of the $bbjj\tau\nu$ final state at the MuC.
@@
  • Jet multiplicity: Exactly 2 light jets and 2 $b$-jets, along with one hadronically decaying tau lepton, with transverse momentum $p_{T}>25$ GeV and pseudorapidity $|\eta|<2.5$ for all QCD jets and the tau jet.
@@
To establish an effective cut-flow, we begin with a detailed analysis of the key kinematic variables that differentiate signal from background. The normalized distributions of the most critical variables are presented in Figure 8: $\Delta R(b_{1},b_{2})$ in the upper-left panel, $\Delta R(j_{1},b_{1})$ in the upper-right panel, ${E_{T}^{\rm miss}}$ in the lower-left panel, and $M(j_{1}j_{2}b_{1}b_{2})$ in the lower-right panel. The signal distributions are depicted as solid lines, while the background distributions are shown as orange histograms. The $\Delta R$ and ${E_{T}^{\rm miss}}$ distributions are presented after applying the basic selection criteria, for the representative case of $m_{H^{\pm}}=130$ GeV. For the $M(j_{1}j_{2}b_{1}b_{2})$ distribution, we show the results for $M_{H^{\pm}}=130{\;{\rm GeV}}$ (blue), 150 GeV (green), and 170 GeV (red) after imposing $\Delta R(b_{1},b_{2})<0.6$, $\Delta R(j_{i},b_{i^{\prime}})|_{i,i^{\prime}=1,2}<0.6$, and ${E_{T}^{\rm miss}}>500{\;{\rm GeV}}$.
@@
Similarly, the angular separation between the leading light jet and the leading $b$-jet, $\Delta R(j_{1},b_{1})$, exhibits a distinct pattern for the signal compared to the background. The signal distribution shows a prominent primary peak at $\Delta R(j_{1},b_{1})\simeq 0.1$, reflecting the fact that $j_{1}$ and $b_{1}$ originate from the same $H^{\pm}$ decay. This feature is consistently observed in other combinations, such as $\Delta R(j_{1},b_{2})$, $\Delta R(j_{2},b_{1})$, and $\Delta R(j_{2},b_{2})$, as expected from the kinematics of the signal process. Unexpectedly, a secondary peak is observed at higher values around $\Delta R\sim 3$ in the signal distribution. We found that it arises from occasional misidentification between light jets and tau-jets. Despite the presence of this small secondary peak, a loose cut of $\Delta R(j_{i},b_{i^{\prime}})<0.6$ for $i,i^{\prime}=1,2$ is still effective in enhancing the signal over the background by retaining the prominent primary peak region.
Another crucial discriminating variable is the missing transverse energy, shown in the lower-left panel of Figure 8. The signal exhibits a significantly higher ${E_{T}^{\rm miss}}$ distribution compared to the background, attributed to the different production mechanisms and mother particles of the neutrino. In the signal process, the $[\tau\nu]$ system originates from the decay of a charged Higgs boson produced through a $2\rightarrow 2$ scattering process at the MuC. The large energy transfer to the $H^{\pm}$ results in a highly energetic neutrino from its subsequent decay, leading to substantial missing transverse energy in the final state. In contrast, the background process involves a $W^{\pm}$ boson produced through a $2\rightarrow 4$ scattering process, which inherently results in a softer energy transfer to the $W^{\pm}$ and, consequently, a less energetic neutrino from its decay. As a result, a stringent cut of ${E_{T}^{\rm miss}}>500$ GeV is highly effective in enhancing the signal significance.
@@
Table 3: Cut-flow for the signal $\mu^{+}\mu^{-}\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$ with $M_{H^{\pm}}$ = 130, 150, and 170 GeV at the 3 TeV MuC. We set $M_{H}=M_{A}=200{\;{\rm GeV}}$, $\tan\beta=10$, and $m_{12}^{2}=3.84\times 10^{3}{\;{\rm GeV}}^{2}$. The significance $\mathcal{S}^{10\%}_{1{\rm ab}^{-1}}$ is calculated considering a 10% background uncertainty and an integrated luminosity of 1 ab-1.
Based on the observations from the key kinematic variable distributions, we propose an effective cut-flow strategy to achieve the high discovery potential at the 3 TeV MuC, as summarized in Table 3. Although we do not present the cut-flow for the negligible ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}/\mu_{\rm f}^{+}\mu_{\rm f}^{-}$, their contributions are included in calculating the significance. In the following discussion, the selection efficiency of each specific cut is measured relative to the preceding cut in the cut-flow.
For the signal process, we set $M_{H}=M_{A}=200{\;{\rm GeV}}$, $\tan\beta=10$, and $m_{12}^{2}=3.84\times 10^{3}{\;{\rm GeV}}^{2}$, and consider three benchmark scenarios with charged Higgs boson masses of $M_{H^{\pm}}=130$, 150, and 170 GeV. After applying the basic selection criteria, the signal cross sections show similar values across the three $M_{H^{\pm}}$ cases, with a slight decrease as $M_{H^{\pm}}$ increases.
The first effective kinematic cut is on the missing transverse energy, with a threshold of ${E_{T}^{\rm miss}}>500$ GeV. This cut results in approximately 70% selection efficiency for the signal and only about 30% for the background, providing an initial suppression of the background while retaining a substantial fraction of the signal events.
The most crucial cut is on the angular distance between the leading and subleading $b$-jets, $\Delta R(b_{1},b_{2})<0.6$. Among the events that satisfy the ${E_{T}^{\rm miss}}$ cut, almost all the signal events survive under this $\Delta R(b_{1},b_{2})$ cut, while only about 30% of the background events remain. At this stage, we already achieve substantial signal significances: $\mathcal{S}\approx 6$ for $m_{H^{\pm}}=130$ GeV and $150$ GeV, and $\mathcal{S}\approx 2.5$ for $m_{H^{\pm}}=170$ GeV. The smaller significance for $m_{H^{\pm}}=170$ GeV is attributed to the reduced production cross section due to the heavier charged Higgs boson mass.
@@
Finally, we impose a condition on the invariant mass of the two $b$-jets and two light jets, such that $M(b_{1}b_{2}j_{1}j_{2})<173$ GeV. Note that we apply an upper bound on $M(b_{1}b_{2}j_{1}j_{2})$ rather than a mass window. This approach is designed to retain as many signal events as possible, since the signal cross sections after the final selection are only of the order of $\mathcal{O}(10)$ ab, resulting in dozens of signal events with a total integrated luminosity of $1$ ab-1. Additionally, the relatively broad resonance peak, due to smearing effects in reconstructing the two $b$-jets and two light jets, is another reason for not imposing a narrow mass window (see Figure 8). This final selection ensures a significant discovery potential for all three benchmark cases. For $M_{H^{\pm}}=130$, 150, and 170 GeV, the final signal significances are 13.7, 13.5, and 6.06, respectively.
To investigate the impacts of higher collision energy on probing the $H^{\pm}\rightarrow t^{*}b$ mode, we consider the 10 TeV MuC with a total integrated luminosity of $10{\;{\rm ab}^{-1}}$. One might naively expect that the 10 TeV MuC would yield higher signal sensitivity than the 3 TeV MuC, as the 10 TeV MuC benefits from the VBS processes of $\mu^{+}\mu^{-}\rightarrow H^{+}H^{-}\nu\bar{\nu}/\mu_{\rm f}^{+}\mu_{\rm f}^{-}$, which have parton-level cross sections more than two times larger than that of the Drell-Yan process (see Table 2).
@@
Table 4: Cut-flow for the signal $\mu^{+}\mu^{-}\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$ at the 10 TeV MuC with the total integrated luminosity of 10 ab-1. For $M_{H^{\pm}}=130{\;{\rm GeV}}$, we present the detailed cut-flow of the cross section of the Drell-Yan process ($\sigma_{\rm sg}^{\rm DY}$), of ${\mu^{+}\mu^{-}}\rightarrow H^{+}H^{-}\nu\bar{\nu}$ ($\sigma_{\rm sg}^{\nu\bar{\nu}}$), and of ${\mu^{+}\mu^{-}}\rightarrow H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ ($\sigma_{\rm sg}^{\mu_{\rm f}\mu_{\rm f}}$). For $M_{H^{\pm}}=170{\;{\rm GeV}}$, we present the cut-flow of the dominant $\sigma_{\rm sg}^{\rm DY}$. The cut of $\Delta R(j_{i},b_{i^{\prime}})<0.2$ collectively denotes four combinations with $i,i^{\prime}=1,2$.
In Table 4, we present the cut-flow of the cross sections for $M_{H^{\pm}}=130{\;{\rm GeV}}$ and $M_{H^{\pm}}=170{\;{\rm GeV}}$ at the 10 TeV MuC. The results for $M_{H^{\pm}}=150{\;{\rm GeV}}$ are omitted as they are very similar to those for $M_{H^{\pm}}=130{\;{\rm GeV}}$, with differences below 1%. The significances are calculated for a total integrated luminosity of $10{\;{\rm ab}^{-1}}$ and a 10% background uncertainty. For $M_{H^{\pm}}=130{\;{\rm GeV}}$, we present the cut-flow for all three production channels: the Drell-Yan process, ${\mu^{+}\mu^{-}}\rightarrow H^{+}H^{-}\nu\bar{\nu}$, and ${\mu^{+}\mu^{-}}\rightarrow H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$, with their respective cross sections denoted by $\sigma_{\rm sg}^{\rm DY}$, $\sigma_{\rm sg}^{\nu\bar{\nu}}$, and $\sigma_{\rm sg}^{\mu_{\rm f}\mu_{\rm f}}$. After the basic selection, the cross sections for the three processes are of the same order of magnitude. However, the ${E_{T}^{\rm miss}}>1{\;{\rm TeV}}$ cut substantially suppresses $\sigma_{\rm sg}^{\nu\bar{\nu}}$ and $\sigma_{\rm sg}^{\mu_{\rm f}\mu_{\rm f}}$, leaving less than 10% of events. In contrast, the Drell-Yan process maintains a high selection efficiency, around 83%. In the case of $M_{H^{\pm}}=170\text{ GeV}$, we present only the Drell-Yan cross sections since the non-Drell-Yan signal processes are negligible after the final selection. However, the signal significances are calculated based on the total cross section from all three processes.
@@
V Conclusions

### v2

@@ 4 $H^{\pm}\rightarrow t^{*}b$ at a Multi-TeV Muon Collider
The Drell-Yan process shows nearly constant cross sections for different $M_{H^{\pm}}$ values at a given $\sqrt{s}$: approximately 3 fb at $\sqrt{s}=3{\;{\rm TeV}}$ and 0.3 fb at $\sqrt{s}=10{\;{\rm TeV}}$. This stability arises because the chosen $M_{H^{\pm}}$ values (see Equation 41) are relatively small compared to the multi-TeV collision energies. The significant decrease in cross sections from $\sqrt{s}=3{\;{\rm TeV}}$ to $\sqrt{s}=10{\;{\rm TeV}}$ reflects the typical behavior of Drell-Yan processes, where cross sections scale as $1/s$. In contrast, the cross sections for ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ moderately increase with $\sqrt{s}$, following the characteristic $\log^{2}({s}/{m_{V}^{2}})$ dependence of the cross sections for VBS processes Han:2021udl .
@@
Next, we present the signal-to-background analysis at the detector level. [^7] In our analysis, we did not include initial state radiation (ISR) or beamstrahlung effects, as these are significantly less pronounced at a muon collider compared to an $e^{+}e^{-}$ collider. The much larger muon mass suppresses both ISR and beamstrahlung, making such corrections less critical for our study. We conducted showering with Pythia version 8.307. A rapid detector simulation was performed using Delphes version 3.5.0, utilizing the delphes_card_
MuonColliderDet.tcl
card. The MuC Delphes card accommodates slightly different values for the $\tau$ tagging and mistagging rates compared to those for the HL-LHC: for $p_{T}\geq 10{\;{\rm GeV}}$, $P_{\tau\to\tau}\simeq 80\%$, $P_{e\rightarrow\tau}\simeq 0.1\%$, and $P_{j\to\tau}\simeq 2\%$.
For jet clustering, we employed the exclusive Valencia algorithm Boronat:2014hva ; Boronat:2016tgd , as implemented in FastJet. The Valencia algorithm is particularly effective at high-energy lepton colliders because it efficiently handles initial state radiation and beam-induced backgrounds (BIBs) by explicitly incorporating beam jets.
We specifically chose the exclusive clustering mode because our signal ($H^{+}H^{-}\rightarrow t^{*}b\tau\nu$) produces a fixed number of jets—exactly five. The exclusive mode halts clustering once the predetermined jet multiplicity is reached, ensuring optimal performance for signals with a known jet count. In contrast, the inclusive clustering algorithm tends to misidentify or merge closely spaced signal jets, especially at multi-TeV muon colliders, where jets from the same parent particle ($H^{\pm}$) become highly collimated.
The exclusive Valencia algorithm has three configurable parameters: the jet radius $R$, and the energy and angular weighting parameters $\beta$ and $\gamma$ that enter the clustering distance measures Boronat:2016tgd . Specifically, $\beta$ and $\gamma$ define the inter-particle distance $d_{ij}$ and the beam distance $d_{iB}$ as:
$d_{ij}$ $=\frac{2\,\text{min}(E_{i}^{2\beta},E_{j}^{2\beta})(1-\cos\theta_{ij})}{R^{2}},$ (42)
$d_{iB}$ $=E_{i}^{2\beta}\sin^{2\gamma}\theta_{iB},$ (43)
where $\theta_{iB}$ is the angle of particle $i$ relative to the beam axis. For robust jet reconstruction, we adopt the canonical settings for the algorithm: $R=0.2$ and $\beta=\gamma=1$.
@@
To identify relevant SM backgrounds for the $bbjj\tau\nu$ final state, we note that the $\tau\nu$ system must originate from a $W^{\pm}$ boson at the MuC. Since $W^{\pm}$ bosons are produced in pairs at the MuC, the two light jets in the final state must arise from the decay of the second $W^{\pm}$. This makes $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}b\bar{b}$ the primary background. Notably, this process also includes top quark pair production as a subset.
There are two additional backgrounds worth considering. The first is $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}jj$, where light quark jets are misidentified as $b$ jets. This includes contributions from $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}Z$ with $Z\to jj$. The second is $\mu^{+}\mu^{-}\rightarrow t\bar{t}Z$, followed by $Z\to\nu\bar{\nu}$.
However, both of these additional backgrounds are negligible compared to the dominant $W^{+}W^{-}b\bar{b}$ process. For the $W^{+}W^{-}jj$ background, the contribution is strongly suppressed by realistic $b$-jet mistagging rates. Specifically, the cross section for $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}jj\rightarrow\tau\nu jjbb$, with kinematic cuts of $p_{T}>20~{}{\;{\rm GeV}}$, $|\eta|<2.5$, and $\Delta R_{ij}>0.4$, is approximately $1.9{\;{\rm ab}}$ at the 3 TeV MuC and $0.4{\;{\rm ab}}$ at the 10 TeV MuC. For the $t\bar{t}Z\to bbjj\tau\nu\nu\bar{\nu}$ background, the corresponding cross sections under the same kinematic cuts are approximately $5.3{\;{\rm ab}}$ at 3 TeV and $0.027{\;{\rm ab}}$ at 10 TeV. In contrast, the cross sections for $\mu^{+}\mu^{-}\to W^{+}W^{-}b\bar{b}$ after the same selection are $549{\;{\rm ab}}$ at 3 TeV and $68.9{\;{\rm ab}}$ at 10 TeV. Given this large disparity, we identify $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}b\bar{b}$ as the dominant background for the $bbjj\tau\nu$ final state at the MuC.
@@
Brief comments on BIBs are in order here. Beam-induced backgrounds arise from the decay of muons in the beam, which produces electrons and positrons. As these interact with machine components, various secondary particles such as photons, electron-positron pairs, hadrons, and neutrinos are generated. Most of these secondary particles result from forward scattering, causing them to be directed along the beam and to have low energies, typically below 1 GeV Bartosik:2024ulr . Given that our final state requires five hard jets in the central region, BIBs are of negligible concern. Thus, we do not consider BIBs a significant background source for our analysis of the $bbjj\tau\nu$ final state at the MuC.
@@
  • Jet multiplicity: Exactly 2 light jets and 2 $b$-jets, along with one hadronically decaying tau lepton, with transverse momentum $p_{T}>25$ GeV and pseudorapidity $|\eta|<2.5$.
@@
To establish an effective cut-flow, we begin with a detailed analysis of the key kinematic variables that differentiate signal from background. The normalized distributions of the critical variables are presented in Figure 8: $\Delta R(b_{1},b_{2})$ in the upper-left panel, $\Delta R(j_{1},b_{1})$ in the upper-right panel, ${E_{T}^{\rm miss}}$ in the lower-left panel, and $M(j_{1}j_{2}b_{1}b_{2})$ in the lower-right panel. The signal distributions are depicted as solid lines, while the background distributions are shown as orange histograms. The $\Delta R$ and ${E_{T}^{\rm miss}}$ distributions are presented after applying the basic selection criteria, for the representative case of $M_{H^{\pm}}=130$ GeV. [^8] The $\Delta R$ and ${E_{T}^{\rm miss}}$ distributions remain almost identical for $M_{H^{\pm}}=130,~{}150,~{}170{\;{\rm GeV}}$ since these masses are much smaller than $\sqrt{s}$ at the multi-TeV MuC. For the $M(j_{1}j_{2}b_{1}b_{2})$ distribution, we show the results for $M_{H^{\pm}}=130{\;{\rm GeV}}$ (blue), 150 GeV (green), and 170 GeV (red) after imposing
$\Delta R(b_{1},b_{2})<0.6,\;\Delta R(j_{i},b_{i^{\prime}})\big{|}_{i,i^{\prime}=1,2}<0.6,$ (44)
${E_{T}^{\rm miss}}>500{\;{\rm GeV}}.$
@@
Similarly, the angular separation between the leading light jet and the leading $b$-jet, $\Delta R(j_{1},b_{1})$, exhibits a distinct pattern for the signal compared to the background. The signal distribution features a prominent primary peak at $\Delta R(j_{1},b_{1})\simeq 0.1$, indicating that $j_{1}$ and $b_{1}$ originate from the same $H^{\pm}$ decay. This trend is consistently observed in other combinations, such as $\Delta R(j_{1},b_{2})$, $\Delta R(j_{2},b_{1})$, and $\Delta R(j_{2},b_{2})$, as expected from the kinematics of the signal process.
Interestingly, the signal distribution also exhibits a secondary peak at larger values around $\Delta R\sim 3$. We find that this arises from occasional misidentification between light jets and tau-jets, which is accounted for in our detector simulation. Specifically, the Delphes setup incorporates realistic mistagging probabilities, including a $P_{j\to\tau}\simeq 2\%$ probability for a light jet to be misidentified as a tau jet, leading to this feature in the distribution.
Another crucial discriminating variable is the missing transverse energy ${E_{T}^{\rm miss}}$, shown in the lower-left panel of Figure 8. The signal exhibits a significantly higher ${E_{T}^{\rm miss}}$ distribution compared to the background, attributed to the different production mechanisms and mother particles of the neutrino. In the signal process, the $[\tau\nu]$ system originates from the decay of a charged Higgs boson produced through a $2\rightarrow 2$ scattering process at the MuC. The large energy transfer to the $H^{\pm}$ results in a more energetic neutrino from its subsequent decay, leading to substantial missing transverse energy in the final state. In contrast, the background process involves a $W^{\pm}$ boson produced through a $2\rightarrow 4$ scattering process, which inherently results in a softer energy transfer to the $W^{\pm}$ and, consequently, a less energetic neutrino from its decay. As a result, a stringent cut of ${E_{T}^{\rm miss}}>500$ GeV is highly effective in enhancing the signal significance.
@@
Table 3: Cut-flow for the signal process $\mu^{+}\mu^{-}\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$ with $M_{H^{\pm}}=130$, 150, and 170 GeV at the 3 TeV MuC. We set $M_{H}=M_{A}=200$ GeV, $\tan\beta=10$, and $m_{12}^{2}=3.84\times 10^{3}$ GeV2. The dominant background process is $\mu^{+}\mu^{-}\to W^{+}W^{-}b\bar{b}$. The significance $\mathcal{S}^{10\%}_{1{\rm ab}^{-1}}$ is calculated assuming a 10% background uncertainty and an integrated luminosity of 1 ab-1.
Based on the observations from the key kinematic variable distributions, we propose an effective cut-flow strategy to achieve the high discovery potential at the 3 TeV MuC, as summarized in Table 3. The significances are calculated for a total integrated luminosity of $1{\;{\rm ab}^{-1}}$ and a 10% background uncertainty. Although we do not present the cut-flow for the negligible ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$, their contributions are included in calculating the significance. In the following discussion, the selection efficiency of each specific cut is measured relative to the preceding cut in the cut-flow.
After applying the basic selection criteria, the signal cross sections show similar values across the three $M_{H^{\pm}}$ cases of 130, 150, and 170 GeV, with a slight decrease as $M_{H^{\pm}}$ increases. The first effective kinematic cut is on the missing transverse energy, requiring ${E_{T}^{\rm miss}}>500$ GeV. This cut achieves approximately 70% selection efficiency for the signal while retaining only about 30% of the background, providing substantial initial background suppression.
The most crucial cut is on the angular distance between the leading and subleading $b$-jets, $\Delta R(b_{1},b_{2})<0.6$. Among events that satisfy the ${E_{T}^{\rm miss}}$ cut, almost all signal events survive this $\Delta R(b_{1},b_{2})$ cut, while only about 30% of background events remain. At this stage, we achieve substantial signal significances: $\mathcal{S}^{10\%}_{1{\rm ab}^{-1}}\approx 6$ for $m_{H^{\pm}}=130$ GeV and $150$ GeV, and $\mathcal{S}^{10\%}_{1{\rm ab}^{-1}}\approx 2.5$ for $m_{H^{\pm}}=170$ GeV. The smaller significance for $m_{H^{\pm}}=170$ GeV is due to the reduced production cross section at higher charged Higgs boson mass.
@@
Finally, we impose a condition on the invariant mass of the two $b$-jets and two light jets, such that $M(b_{1}b_{2}j_{1}j_{2})<173$ GeV. We apply an upper bound on $M(b_{1}b_{2}j_{1}j_{2})$ rather than a mass window. This approach is designed to retain as many signal events as possible, since the signal cross sections after the final selection are only of the order of $\mathcal{O}(10)$ ab, resulting in dozens of signal events with a total integrated luminosity of $1$ ab-1. Additionally, the relatively broad resonance peak, due to smearing effects in reconstructing the two $b$-jets and two light jets, is another reason for not imposing a narrow mass window (see Figure 8). This final selection ensures a significant discovery potential for all three benchmark cases. For $M_{H^{\pm}}=130$, 150, and 170 GeV, the final signal significances are 13.7, 13.5, and 6.06, respectively.
To explore how higher collision energy affects the sensitivity to the $H^{\pm}\rightarrow t^{*}b$ mode, we analyze the 10 TeV MuC with a total integrated luminosity of $10{\;{\rm ab}^{-1}}$. One might expect enhanced signal sensitivity compared to the 3 TeV MuC, given both the larger luminosity and the increased importance of VBS processes ($\mu^{+}\mu^{-}\rightarrow H^{+}H^{-}\nu\bar{\nu}/\mu_{\rm f}^{+}\mu_{\rm f}^{-}$): see Table 2.
@@
Table 4: Cut-flow for the signal process $\mu^{+}\mu^{-}\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$ at the 10 TeV MuC with a total integrated luminosity of 10 ab-1. The dominant background process is $\mu^{+}\mu^{-}\to W^{+}W^{-}b\bar{b}$. For $M_{H^{\pm}}=130$ GeV, we present the detailed cut-flow of the cross section for the Drell-Yan process ($\sigma_{\rm sg}^{\rm DY}$), ${\mu^{+}\mu^{-}}\rightarrow H^{+}H^{-}\nu\bar{\nu}$ ($\sigma_{\rm sg}^{\nu\bar{\nu}}$), and ${\mu^{+}\mu^{-}}\rightarrow H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ ($\sigma_{\rm sg}^{\mu_{\rm f}\mu_{\rm f}}$). For $M_{H^{\pm}}=170$ GeV, we present the cut-flow of the dominant $\sigma_{\rm sg}^{\rm DY}$. The cut $\Delta R(j_{i},b_{i^{\prime}})<0.2$ collectively denotes four possible combinations with $i,i^{\prime}=1,2$.
In Table 4, we present the cut-flow of the cross sections for $M_{H^{\pm}}=130{\;{\rm GeV}}$ and $M_{H^{\pm}}=170{\;{\rm GeV}}$ at the 10 TeV MuC. The results for $M_{H^{\pm}}=150{\;{\rm GeV}}$ are omitted as they are very similar to those for $M_{H^{\pm}}=130{\;{\rm GeV}}$, with differences below 1%. The significances are calculated for a total integrated luminosity of $10{\;{\rm ab}^{-1}}$ and a 10% background uncertainty. For $M_{H^{\pm}}=130{\;{\rm GeV}}$, we present the cut-flow for all three production channels: the Drell-Yan process, ${\mu^{+}\mu^{-}}\rightarrow H^{+}H^{-}\nu\bar{\nu}$, and ${\mu^{+}\mu^{-}}\rightarrow H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$, with their respective cross sections denoted by $\sigma_{\rm sg}^{\rm DY}$, $\sigma_{\rm sg}^{\nu\bar{\nu}}$, and $\sigma_{\rm sg}^{\mu_{\rm f}\mu_{\rm f}}$. After the basic selection, the cross sections for the three processes are of the same order of magnitude. However, the ${E_{T}^{\rm miss}}>1{\;{\rm TeV}}$ cut substantially suppresses $\sigma_{\rm sg}^{\nu\bar{\nu}}$ and $\sigma_{\rm sg}^{\mu_{\rm f}\mu_{\rm f}}$, leaving less than 10% of events. In contrast, the Drell-Yan process maintains a high selection efficiency, around 83%. In the case of $M_{H^{\pm}}=170\text{ GeV}$, we present only the Drell-Yan cross sections since the other processes are negligible after the final selection. However, the signal significances are calculated based on the total cross section from all three processes.
@@
A discussion of CLIC’s discovery potential for the $H^{\pm}\to t^{*}b$ mode merits attention. In its final stage, CLIC Brunner:2022usy will operate at a c.m. energy of 3 TeV with an integrated luminosity of approximately $5{\;{\rm ab}^{-1}}$. Unlike a multi-TeV MuC, where detection is limited to $|\eta|<2.5$ due to BIBs, CLIC faces no such restrictions as an electron-positron collider. As a result, a 3 TeV CLIC is expected to have superior sensitivity to the $H^{\pm}\to t^{*}b$ mode compared to a 3 TeV MuC. The intermediate stage at $\sqrt{s}=1.5{\;{\rm TeV}}$ should also provide substantial discovery reach. Given the distinct collider characteristics, a dedicated study of this channel at CLIC would be valuable for future study.
5 Conclusions